# Mathematical modelling of the vitamin C clock reaction: a study of two kinetic regimes

### v1

@@ 2 Model formulation
Introducing the variables $P(t)$ for hydrogen peroxide and $D(t)$ for hypoiodous acid, these reactions are simplified as,
@@
Figure 1: Comparison of numerical and asymptotic solutions (found in sections 3 and 4) to the dimensionless problem (12) against logarithmic time. Parameter values (chosen arbitrarily): $\epsilon=10^{-2}$, $\beta=0.6$, $\gamma=0.7$, $\sigma=0.8$, $\phi=0.2$, with (a) $\rho=2$ (M-HP), and (b) $\rho=\hat{\rho}\epsilon^{-1}=90$ (H-HP).
@@ 3.3 Region II: induction period
A similar equation to (32) was given in ref. [10] through a quasi-steady argument, although not including the $\phi n_{0}$ term in the denominator. Equation (32) is one of the two formulae that will be tested experimentally in this paper.
@@ 3.5 Region IV: equilibriation
A comparison between numerical and leading order asymptotic solutions in each region in dimensionless variables is given in figure 1(a), for a specific set of parameters, with small parameter chosen as $\epsilon=10^{-2}$. In the absence of detailed measurement of the reaction rates in the system, this choice is arbitrary. As shown in figure 1(a), there is a small but noticeable difference between the asymptotic and numerical solutions as $C(t)$ approaches zero in the induction region which results in a small discrepancy in switchover time (figure 2(a)). By plotting as a function of $\epsilon$ the relative difference in the asymptotic switchover time value (equation (31)) and the point at which the numerical solution falls below a threshold value of $\epsilon^{-1}$ (figure 2(b)), this discrepancy converges to zero. As would be expected from taking only the leading order terms of the asymptotic expansion, the convergence is approximately linear in $\epsilon$.
(a) (b)
Refer to caption Refer to caption
@@ 4.4 Region IV: equilibriation
Figure 1(b) compares the high hydrogen peroxide asymptotic solutions in each region with a numerical solution computed with an arbitrary set of parameters (details in caption) corresponding. The solutions are quite close in their region of validity, particularly for the trajectories of $P$, $I$ and $C$, although as time progresses, a significant drift develops between the numerically computed value of $P$ and the constant approximation produced by the asymptotic analysis; the discrepancy appears to be on the order of $\epsilon^{1/2}$. This discrepancy could be reduced by seeking the next order term in the asymptotic expansion, although this quantity is less critical to determine than the clock chemicals $I$ and $C$ which directly determine the switchover time.
@@ 5.4 Results
Figure 4(a–d) shows the outcome of the fitting series experiments $N_{M}$, $C_{M}$, $N_{H}$ and $C_{H}$. The quality of the fit appears excellent, with the maximum relative error rarely exceeding 10%. The estimated parameter values of $\hat{\phi}=0.158$ (95% confidence interval $[0.152,0.161]$) and $\hat{k}_{2}=0.0663$ (95% CI $[0.639,0.693]$ are shown in figure 4(e) along with bootstrap sample distributions. The testing series $H_{1}$ and $H_{2}$ using the estimated parameters provide at least as close a match between experiment and the fitted model (figure 5) as for the fitted data, providing confidence in the out-of-sample applicability of the model and accuracy of the fitted parameters.
(a) (b)
Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption
@@ 6 Discussion
The second asymptotic analysis of the paper concerned the high hydrogen peroxide case already considered by Kerr et al. Heuristically one can argue that when hydrogen peroxide levels are sufficiently high, the rate-limiting step is no longer such, and so the overall reaction $2I^{-}\rightarrow I_{2}$ can be modelled by the quadratic kinetics that would follow from the law of mass action, and hydrogen peroxide levels need not be modelled explicitly as they are not significantly reduced. The analysis of section 4 put this informal argument on a more quantitative basis through elucidating how a hydrogen peroxide concentration of order $\epsilon^{-1}$ provides the balance described above. Similarly to Kerr et al. we found four asymptotic regions corresponding to (I) initial reduction of vitamin C due to non-zero initial iodine, which is simultaneous with iodide, iodine and hypoiodous acid reaching quasi-equilibrium, (II) induction, (III) crossover and (IV) evolution to final equilibrium. The leading order solution for region IV could not be solved exactly, however a linearisation about the equilibrium combined with matching to region II enabled a relatively accurate approximate solution to be found; however the leading order analysis was not even in principle able to capture the significant consumption of hydrogen peroxide seen in region IV. The region II solution enabled the deduction of a ‘full’ switchover time formula (71), along with a simplified version that could be justified rigorously under certain assumptions regarding parameters (75) and a highly simplified heuristic version for comparison with Kerr et al. (76). Due to involving only two free parameters, it was the latter formula that we subsequently used for fitting experimental data. The full formula was found to converge to the numerical solution as $\epsilon\rightarrow 0$ sublinearly, suggesting that the next order correction may be sublinear.
@@
Data accessibility: data are available at the repository [23].

### v2

@@ 2 Model formulation
Introducing the variables $P(t)$ for hydrogen peroxide and $Q(t)$ for hypoiodous acid, these reactions are simplified as,
@@
Figure 1: Comparison of numerical solutions (solid lines) and asymptotic solutions (dotted lines), as found in sections 3 and 4 respectively to the dimensionless problem (12) plotted against logarithmic time. Parameter values (chosen arbitrarily): $\epsilon=10^{-2}$, $\beta=0.6$, $\gamma=0.7$, $\sigma=0.8$, $\phi=0.2$. (a) Moderate hydrogen peroxide case M-HP with $\rho=2$; (b) high hydrogen peroxide case H-HP with $\hat{\rho}=0.9$ so that $\rho=\hat{\rho}\epsilon^{-1}=90$.
@@ 3.3 Region II: induction period
A similar equation to (32) was given in ref. [10] through a quasi-steady argument and making the phenomenological assumption that the slow reaction is linear in iodide concentration, (although not including the $\phi n_{0}$ term in the denominator). However, by accounting for the sub-steps of the slow reaction, the dynamics of this regime emerge naturally from the law of mass action. Equation (32) is one of the two formulae that will be tested experimentally in this paper.
@@ 3.5 Region IV: equilibration
A comparison between numerical and leading order asymptotic solutions in each region in dimensionless variables is given in figure 1(a), for a specific set of parameters, with small parameter chosen as $\epsilon=10^{-2}$. In the absence of detailed measurement of the reaction rates in the system, this choice is arbitrary. Additional detail on the dynamics for $Q(t)$ in region Ia is given in Appendix A, figure 6. As shown in figure 1(a), there is a small but noticeable difference between the asymptotic and numerical solutions as $C(t)$ approaches zero in the induction region which results in a small discrepancy in switchover time (figure 2(a)). By plotting as a function of $\epsilon$ the relative difference in the asymptotic switchover time value (equation (31)) and the point at which the numerical solution falls below a threshold value of $\epsilon^{-1}$ (figure 2(b)), this discrepancy converges to zero. As would be expected from taking only the leading order terms of the asymptotic expansion, the convergence is approximately linear in $\epsilon$.
@@ 4.2 Region II: Induction period
Because equation (75) includes an inverse dependence on $p_{0}$, the question may be posed (see open peer review accompanying this article) whether this dependence predicts that switchover time can be made arbitrarily small in practice simply by increasing $p_{0}$? However, sufficiently large values of $p_{0}$ would correspond to $p_{0}/c_{0}$ no longer being merely $O(\epsilon^{-1})$, which would be outside of the range of validity of the present analysis by changing the relative orders of magnitudes of the reaction steps. The implications of taking $p_{0}/c_{0}$ at the next order of magnitude $O(\epsilon^{-2})$ are discussed briefly in Appendix B; in short, the analysis predicts that the switchover time approaches a non-zero value which has no quantitative dependence on $p_{0}$.
@@ 4.4 Region IV: equilibration
Figure 1(b) compares the high hydrogen peroxide asymptotic solutions in each region with a numerical solution computed with an arbitrary set of parameters (details in caption). The solutions are quite close in their region of validity, particularly for the trajectories of $P$, $I$ and $C$, although as time progresses, a significant drift develops between the numerically computed value of $P$ and the constant approximation produced by the asymptotic analysis; the discrepancy appears to be on the order of $\epsilon^{1/2}$. This discrepancy could be reduced by seeking the next order term in the asymptotic expansion, although this quantity is less critical to determine than the clock chemicals $I$ and $C$ which directly determine the switchover time.
@@ 5.4 Results
Figure 4(a–d) shows the outcome of the fitting series experiments $N_{M}$, $C_{M}$, $N_{H}$ and $C_{H}$. Qualitatively, the results show that in both low and high hydrogen peroxide regimes, switchover time declines as initial iodine concentration is increased, and increases nearly linearly with initial vitamin C concentration – fundamentally, as substrate is increased, inhibitor is used up more quickly, whereas as inhibitor is increased, it lasts longer. The quality of the fit appears excellent, with the maximum relative error rarely exceeding 10%. The estimated parameter values of $\hat{\phi}=0.158$ (95% confidence interval $[0.152,0.161]$) and $\hat{k}_{2}=0.0663$ (95% CI $[0.639,0.693]$ are shown in figure 4(e) along with bootstrap sample distributions.
The testing series $H_{1}$ and $H_{2}$ using the estimated parameters provide at least as close a match between experiment and the fitted model (figure 5) as for the fitted data, providing confidence in the out-of-sample applicability of the model and accuracy of the fitted parameters. Qualitatively, the data show switchover time decreasing as hydrogen peroxide concentration is increased, corresponding to the slow reaction being promoted through additional availability of a key reactant in its rate-limiting step.
@@ 6 Discussion
The second asymptotic analysis of the paper concerned the high hydrogen peroxide case already considered by Kerr et al. Heuristically one can argue that when hydrogen peroxide levels are sufficiently high, the rate-limiting step is no longer such, and so the overall reaction $2I^{-}\rightarrow I_{2}$ can be modelled by the quadratic kinetics that would follow from the law of mass action, and hydrogen peroxide levels need not be modelled explicitly as they are not significantly reduced. The analysis of section 4 put this informal argument on a more quantitative basis through elucidating how a hydrogen peroxide concentration of order $\epsilon^{-1}$ provides the balance described above. Similarly to Kerr et al. we found four asymptotic regions corresponding to (I) initial reduction of vitamin C due to non-zero initial iodine, which is simultaneous with iodide, iodine and hypoiodous acid reaching quasi-equilibrium, (II) induction, (III) crossover and (IV) evolution to final equilibrium. The leading order solution for region IV could not be solved exactly, however a linearisation about the equilibrium combined with matching to region II enabled a relatively accurate approximate solution to be found; however the leading order analysis was not even in principle able to capture the significant consumption of hydrogen peroxide seen in region IV.
One of the goals of this paper has been to show how previously-reported formulae for the switchover time can be recovered from our unified model, when considered in the appropriate concentration regime. As previously, the region II solution enabled the deduction of a ‘full’ switchover time formula (71) for the full H-HP regime. This full formula was shown to be equivalent to a simplified version that could be justified rigorously under certain assumptions regarding parameters (75) and a highly simplified heuristic version for comparison with Kerr et al. (76). Due to involving only two free parameters, it was the latter formula that we subsequently used for fitting experimental data. The full formula was found to converge to the numerical solution as $\epsilon\rightarrow 0$ sublinearly, suggesting that the next order correction may be sublinear.
@@
Data accessibility: data are available at the Dryad repository [23].
@@
Appendix A Quasi-equilibrium for $Q(t)$ in the M-HP regime
The establishment of quasi-equilibrium for $Q(t)$ (Figure 1(a), region Ia) is displayed in more detail in figure 6, with vertical axis for concentration and linear axis for time. It is evident that $Q(t)=O(\epsilon)$ and $I(t)=O(\epsilon^{2})$.
Figure 6: Comparison of numerical solutions (solid line) and asymptotic solutions (dotted line), replotted from figure 1(a), region Ia to resolve the dynamics for $Q(t)$.
Appendix B Very high hydrogen peroxide regime
During open peer review, a query was raised as to the behaviour of $t_{sw}$ as $p_{0}$ is made arbitrarily large – specifically, does it follow from equation (75) that $t_{sw}$ should approach zero? However, the results of section 4 rest on the assumption that the initial hydrogen peroxide level is on the order of $\epsilon^{-1}$ relative to vitamin C, i.e. $\rho=\epsilon^{-1}\hat{\rho}$ with $\hat{\rho}$ being order 1. To explore the next order of magnitude in relative hydrogen peroxide concentration, below we briefly consider how the dynamics up to switchover change when instead $\rho=\epsilon^{-2}\hat{\rho}$. The dimensionless equation for $Q$ is then,
$\frac{dQ}{dt}=\hat{\rho}(1-Q-2I)P-\epsilon\gamma\sigma(1-Q-2I)Q-\beta\hat{\rho}QP.$ (93)
Qualitatively, the first and third terms on the right hand side are now order 1, which means that the “$k_{2}$” and “$k_{4}$” reactions predominate, and $Q$ reaches an order 1 quasi-equilibrium in order 1 time. In region I, the leading order behaviour is then,
$\frac{dQ_{0}}{dt}=\hat{\rho}\left(1-(1+\beta)Q_{0}-2I_{0}\right)P_{0},$ (94)
which approaches a steady state $Q_{0}(t)\rightarrow(1+\beta)^{-1}$ as $t\rightarrow\infty$.
Region II then again corresponds to the rescaling $t=\epsilon^{-1}\tau$ and $I=\epsilon\tilde{I}$; the dynamics for $Q(\tau)$ are then,
$\frac{dQ}{d\tau}=\epsilon^{-1}\hat{\rho}\left(1-Q-2\epsilon\tilde{I}\right)P-\gamma\sigma\left(1-Q-2\epsilon\tilde{I}\right)Q-\epsilon^{-1}\beta\hat{\rho}QP,$ (95)
which at leading order yields the constant solution $Q_{0}(\tau)=(1+\beta)^{-1}$. The remaining leading order equations for $P_{0}$, $\tilde{I}_{0}$ and $C_{0}$ are unchanged from equations (61). However with the modified value of $Q_{0}$, the solution for $C_{0}$ then becomes,
$C_{0}(\tau)=1-\sigma\phi-\frac{\beta\gamma\sigma^{2}}{(1+\beta)^{2}}\tau,$ (96)
from which it follows that,
$\tau_{sw}=\frac{(1-\sigma\phi)(1+\beta)^{2})}{\beta\gamma\sigma^{2}}.$ (97)
Returning to dimensional variables, the switchover time formula in this regime is then,
$t_{sw}=\frac{c_{0}-\phi n_{0}}{(k_{3}/k_{1})(k_{4}/k_{2})(1+(k_{4}/k_{2}))^{-2}n_{0}^{2}},$ (98)
which takes a similar form to the formula of Kerr et al. [9]. The switchover time depends on the reaction rates only through the ratios $k_{3}/k_{1}$ (2nd forward step of slow reaction relative to fast reaction) and $k_{4}/k_{2}$ (reverse step of slow reaction relative to first forward step). Qualitatively, it is the second ($k_{3}$) step that is now rate limiting for the slow reaction, with the rate set by the $Q$-level that emerges from the quasi-steady balance of the hydrogen peroxide-dependent reactions ($k_{2}$ and $k_{4}$).
In conclusion, we predict that once the hydrogen peroxide level is sufficiently high, the analysis of section 4 is no longer valid and $t_{sw}$ does not approach zero. Instead, the switchover time formula depends on vitamin C and iodine concentrations similarly to the high hydrogen peroxide case, but the quantitative dependence on hydrogen peroxide level is lost.